Emer is 27 years older than his son. In 10 years he will be twice as old as his son. What is the age of both?

Answer by lynnlo(4176)

E============EMBER
S============SON
E=S+27
(E+10)=2*(S+10)
E=S+27
(E+10)=2*(S+10)
====SIMPLIFY====
E-S=27
E-2*S=2*10-10